# Accelerating Diffusion Models in Offline RL via Reward-Aware Consistency Trajectory Distillation

### v1

@@ Abstract
Although diffusion models have achieved strong results in decision-making tasks, their slow inference speed remains a key limitation. While the consistency model offers a potential solution, its applications to decision-making often struggle with suboptimal demonstrations or rely on complex concurrent training of multiple networks. In this work, we propose a novel approach to consistency distillation for offline reinforcement learning that directly incorporates reward optimization into the distillation process. Our method enables single-step generation while maintaining higher performance and simpler training. Empirical evaluations on the Gym MuJoCo benchmarks and long horizon planning demonstrate that our approach can achieve an $8.7\%$ improvement over previous state-of-the-art while offering up to 142×142$\times$142 × speedup over diffusion counterparts in inference time.
@@ 1 Introduction
To address this limitation, various diffusion acceleration techniques have been proposed, including ordinary or stochastic differential equations (ODE or SDE) solvers with flexible step sizes (Song et al., 2020a; Lu et al., 2022; Karras et al., 2022), sampling step distillation (Song et al., 2023; Kim et al., 2023) and improved noise schedules and parametrizations (Salimans and Ho, 2022; Song and Dhariwal, 2024). In particular, consistency distillation (Song et al., 2023) has emerged as one of the most promising solutions for image generation, in which a many-step diffusion model serves as a teacher to train a student consistency model that achieves comparable performance while enabling faster sampling through a single-step or few-step generation process.
@@
This breakthrough has sparked considerable interest in applying consistency distillation to decision-making tasks. However, current applications either adopt a behavior cloning approach (Lu et al., 2024; Prasad et al., 2024; Wang et al., 2024) or integrate few-step diffusion based samplers in actor-critic frameworks (Chen et al., 2023; Ding and Jin, 2023; Li et al., 2024). While promising, these approaches face inherent challenges: behavior cloning performs well only with expert demonstrations but struggles with suboptimal data (e.g., median-quality replay buffers), while the actor-critic paradigm introduces concurrent training of multiple networks and sensitive hyperparameters, exacerbating training complexity, instability, and computational overhead.
@@
This raises an important question: can we develop a more effective approach to consistency distillation specifically tailored for offline RL? We address this challenge by introducing a novel method that directly incorporates reward optimization into the consistency distillation process. Our approach begins with a pre-trained unconditional diffusion policy and augments the standard consistency trajectory distillation (Kim et al., 2023) with an explicit reward objective. The vanilla consistency trajectory distillation helps the student model cover the diverse behavior patterns learned by the teacher. The additional objective encourages the student consistency model to sample actions that yield higher rewards, effectively steering the model toward selecting optimal trajectories from the multi-modal distributions captured by the teacher diffusion model.
@@
Another key advantage of our method lies in its training simplicity. The reward model can be trained independently from the teacher diffusion model and the distillation process, avoiding the complexity of concurrent multi-network training present in actor-critic methods. Our method also eliminates the need for training noise-aware reward models, unlike existing guided diffusion sampling approaches (Janner et al., 2022). By integrating reward optimization directly into the distillation process, our method achieves superior performance, enables efficient single-step generation, and maintains straightforward training procedures.
@@
We demonstrate the performance and sampling efficiency of our RACTD on the suboptimal heterogenous D4RL Gym-MuJoCo benchmark and challenging long-horizon planning task Maze2d (Fu et al., 2020). Our method demonstrates both superior performance and substantial sampling efficiency compared to existing approaches, achieving an $8.7\%$ improvement compared to existing state-of-the-art (SOTA) and a $142$-fold reduction in sampling time.
@@
Our contributions include: (1) We propose a novel reward-aware consistency distillation method for offline RL that enables single-step generation while achieving superior performance, (2) We demonstrate that our approach enables decoupled training without the complexity of concurrent multi-network optimization or noise aware reward model training, (3) Through comprehensive experiments on multi-modal suboptimal dataset and long-horizon planning tasks, we show that our method achieves $8.7\%$ improvement over prior SOTA while achieving up to 142×142$\times$142 × speedup. We will publicly release our code upon acceptance of this paper.
@@ 2.3 Consistency Trajectory Distillation
Since the distance metric $d$ is defined on the clean data space and may not be well-defined in the noisy data space, in practice we further map all predictions to time 00 using the student model. Formally, denote the student model as $G_{\theta}$, the CTM loss is defined as:
$\small\mathcal{L}_{\mathrm{CTM}}=\mathbb{E}\left[d\left(G_{sg(\theta)}(\hat{\bm{x}}_{k}^{(t)},k,0),G_{sg(\theta)}(\bm{x}_{k}^{(t,u)},k,0)\right)\right]$ (5)
@@
In addition to the CTM loss, consistency trajectory distillation also incorporates the DSM loss to enforce its samples to be close to the training data. The DSM loss for the student model is the same as the one for EDM in Equation 4:
@@
In Figure 2 we provide a visualization of these objectives on a PFODE trajectory. After the distillation, the student model can perform “anytime-to-anytime” jumps along the PFODE trajectory. One-step sampling can then be achieved by calculating $\hat{\bm{x}}^{(T)}_{0}=G_{\theta}(\bm{x}_{T},T,0)$.
@@ 3.1 Motivation and Intuition
Diffusion models and their consistency-based counterparts have demonstrated promising results in decision-making tasks, particularly in capturing multimodal behavior patterns (Janner et al., 2022; Chi et al., 2023; Ajay et al., 2022). Common recipes for using these models in decision-making generally fall into one of the three paradigms: (1) training a diffusion or consistency model via behavior cloning and deploying it directly as a policy (Chi et al., 2023; Ajay et al., 2022); (2) integrating the model into actor-critic frameworks (Wang et al., 2022; Hansen-Estruch et al., 2023; Ding and Jin, 2023); or (3) using the model as a planner through guided diffusion sampling (Janner et al., 2022).
While behavior cloning pipelines perform well when trained on expert demonstrations, they often struggles with suboptimal datasets (e.g., medium-replay buffers) collected from diverse behavior policies. Such datasets typically exhibit complex multimodal behavior patterns, where only some modes lead to high rewards. Although one could potentially use rejection sampling to filter out low-reward training data, this approach becomes prohibitively sample inefficient, particularly as the quality of the training data deteriorates. On the other hand, to generate high reward actions, actor-critic approaches require concurrent training of multiple neural networks with sensitive hyperparameters. Finally, guided diffusion sampling necessitates training noise-aware reward models and multi-step sampling, which could be detrimental for time-sensitive decision-making tasks like self-driving.
So how can we better leverage potentially suboptimal datasets to design a diffusion-based single-step sampling model with simple training procedure? Our key idea is to utilize the multimodal information captured by the teacher model and encourage the student model to favor the high reward modes. We achieve this by incorporating a reward objective directly in the consistency distillation process. Since our student model can achieve single-step denoising, we can incorporate a reward model trained in the clean sample space and avoid the multi-step reward optimization for diffusion models.
@@ 3.2 Modeling Action Sequences
Before introducing our reward-aware consistency trajectory distillation, we would like to first clearly establish the modeling formulation in our method. When applying diffusion models to RL, several modeling choices are available: modeling actions (as a policy), modeling rollouts (as a planner), or modeling state transitions (as a world model). While more comprehensive modeling approaches can offer advantages, particularly in long-horizon tasks, they also introduce additional complexity and computational challenges.
Following Chi et al. (2023), we adopt a balanced approach: modeling a fixed-length sequence of future actions conditioned on a fixed-length sequence of observed states. This formulation ensures consecutive actions form coherent sequences, and reduces the chances of generating idle actions.
[…]
@@ 3.3 Reward-Aware Consistency Trajectory Distillation
Having established the formulation, we now present our approach to integrating reward optimization into the consistency trajectory distillation process.
Let $R_{\psi}$ be a pre-trained differentiable return-to-go network (i.e. reward model) that takes the state $\bm{s}_{n}$ and action $\bm{a}_{n}$ at rollout time $n$ as inputs and predicts the future discounted cumulative reward $\hat{r}_{n}=\sum_{j=0}^{H-n}\gamma^{j}r_{n+j}$. When the student model generates a prediction $\vec{\bm{a}}_{n}\mid\vec{\bm{s}}_{n}=\hat{\bm{x}}^{(T)}_{0}=G_{\theta}(\bm{x}_{T},T,0)$, we extract the action at time $n$, denoted as $\hat{\bm{a}}_{n}$, from the predicted sequence and pass it along with $\bm{s}_{n}$ to the frozen reward model $R_{\psi}$ to estimate $\hat{r}_{n}$. The goal of our reward-aware training is to maximize the estimated discounted cumulative reward. Mathematically, the reward objective is defined as
@@
where $\alpha$, $\beta$, and $\sigma$ are hyperparameters to balance different loss terms.
@@ 3.4 Decoupled Training
A key advantage of our method of combining reinforcement learning, diffusion models, and consistency distillation is the ability to support fully decoupled training of all components. Traditional actor-critic frameworks, which are commonly used to incorporate diffusion models into reinforcement learning, require simultaneous training of multiple neural networks. This concurrent optimization presents considerable challenges, often demanding extensive hyperparameter tuning and careful balancing of different learning objectives.
Guided diffusion sampling, as proposed in Janner et al. (2022), offers an alternative approach by taking inspiration from classifier guided diffusion (Dhariwal and Nichol, 2021; Song et al., 2020b). However, these classifiers (i.e. reward models) require noise-aware training that cannot be separated from the diffusion model. Also, predicting the correct reward from highly corrupted input could be very challenging, which can lead to inaccurate guidance that accumulates during its multi-step sampling process.
@@ 3.5 Reward Objective as Mode Selection
In offline RL, models often have to learn from datasets containing behaviors of varying quality. While diffusion models excel at capturing these diverse behavioral modes, they inherently lack the ability to differentiate between actions that lead to high versus low rewards. Our RACTD addresses this limitation by transforming the reward-agnostic teacher diffusion sampling distribution into one that preferentially samples from high-reward modes, We empirically verify this through a comparative analysis using the D4RL hopper-medium-expert dataset Fu et al. (2020), which contains an equal mixture of expert demonstrations and suboptimal rollouts from a partially trained policy.
@@
Figure 3 illustrates the reward distributions of rollouts sampled from three models: the unconditioned teacher, unconditioned student, and RACTD. The dataset (grey) exhibits two distinct modes corresponding to medium-quality and expert rollouts. The unconditioned teacher model (blue) accurately captures this bimodal distribution, and the unconditioned student model (orange) faithfully replicates it. In contrast, our RACTD (green) concentrates its samples on the higher-reward mode, demonstrating that our reward guidance effectively identifies and selects optimal behaviors from the teacher’s multi-modal distribution. We also include the discussion between sample diversity and mode selection in Appendix D.
@@ 4 Experiment
In this section, we conduct experiments to demonstrate: (1) the effectness of our RACTD, especially when the teacher model is trained with suboptimal data, (2) the capability of accurately capturing the complex and diverse high dimensional behavior patterns with only one denoising step, and (3) the speed-up achieved over the teacher model and existing policy-based diffusion models.
@@ 4.1 Offline RL
We compare our approach against a comprehensive set of baselines, including vanilla behavior cloning (BC) and Consistency Policy (Consistency BC) (Ding and Jin, 2023); model-free RL algorithms CQL (Kumar et al., 2020) and IQL (Kostrikov et al., 2021); model-based algorithms Trajectory Transformer (TT) (Janner et al., 2021), MOPO (Yu et al., 2020), MOReL (Kidambi et al., 2020), MBOP (Argenson and Dulac-Arnold, 2020); autoregressive model Decision Transformer (DT) Chen et al. (2021); diffusion-based planner Diffuser Janner et al. (2022); and diffusion-based actor-critic methods Diffusion QL (Wang et al., 2022) and Consistency AC (Ding and Jin, 2023).
@@
We first evaluate our method and the baselines on D4RL Gym-MuJoCo (Fu et al., 2020), which is a popular offline RL benchmark that contains three environments (hopper, walker2d, halfcheetah) of mixtures of varying quality data (medium-replay, medium, medium-expert).
Following the setups in the prior works, we employ online model selection for evaluation by using the best-performing checkpoint observed during training. Results for non-diffusion-based models and Diffuser are taken from Janner et al. (2022), and results for Diffusion QL and Consistency AC/BC are sourced from their respective papers. The results for RACTD are reported as the mean and standard error over 100 planning seeds. We use $h=1,c=16$ and closed-loop plannign in all experiments.
Table 1: (Offline RL) Performance of RACTD and a variety of baselines on the D4RL Gym-MujoCo benchmark. The best score is emphasized in bold and the second-best is underlined.
Method Medium Expert Medium Medium Replay   Avg   NFE

### v2

# Accelerating Diffusion Planners in Offline RL via Reward-Aware Consistency Trajectory Distillation

@@ Abstract
Although diffusion models have achieved strong results in decision-making tasks, their slow inference speed remains a key limitation. While consistency models offer a potential solution, existing applications to decision-making either struggle with suboptimal demonstrations under behavior cloning or rely on complex concurrent training of multiple networks under the actor-critic framework. In this work, we propose a novel approach to consistency distillation for offline reinforcement learning that directly incorporates reward optimization into the distillation process. Our method achieves single-step sampling while generating higher-reward action trajectories through decoupled training and noise-free reward signals. Empirical evaluations on the Gym MuJoCo, FrankaKitchen, and long horizon planning benchmarks demonstrate that our approach can achieve a $9.7\%$ improvement over previous state-of-the-art while offering up to $142\times$ speedup over diffusion counterparts in inference time.
@@ 1 Introduction
To address this limitation, various diffusion acceleration techniques have been proposed, including ordinary or stochastic differential equations (ODE or SDE) solvers with flexible step sizes (Song et al., 2020a; Lu et al., 2022; Karras et al., 2022), sampling step distillation (Song et al., 2023; Kim et al., 2023) and improved noise schedules and parametrizations (Salimans & Ho, 2022; Song & Dhariwal, 2024). In particular, consistency distillation (Song et al., 2023) and consistency trajectory models (Kim et al., 2023) have emerged as one of the most promising solutions for image generation, in which a many-step diffusion model serves as a teacher to train a student consistency model that achieves comparable performance while enabling faster sampling through a single-step or few-step generation process.
@@
This breakthrough has sparked considerable interest in applying consistency-based distillation to decision-making tasks. However, current applications either adopt a behavior cloning approach (Lu et al., 2024; Prasad et al., 2024; Wang et al., 2024) or integrate few-step diffusion based samplers in actor-critic frameworks (Chen et al., 2023; Ding & Jin, 2023; Li et al., 2024). While promising, these approaches face inherent challenges: behavior cloning performs well only with expert demonstrations but struggles with suboptimal data (e.g., median-quality replay buffers), while actor-critic methods require concurrent optimization of multiple networks from scratch with sensitive hyperparameters, leading to training complexity, instability, and computational overhead.
@@
This raises an important question: can we develop a more effective approach to consistency-based distillation that tackles the two fundamental challenges introduced in previous framework: suboptimal data and the concurrent training of actor critic networks for offline RL? We address this challenge by introducing a novel method that directly incorporates reward optimization into the consistency trajectory distillation process. Our approach builds on existing infrastructure, leveraging a pre-trained diffusion planner as teacher and a standalone reward model to augment the standard consistency trajectory distillation (Kim et al., 2023) with an explicit reward objective for a single-step sampling student. The vanilla consistency trajectory distillation helps the student planner distill the diverse behavior modes learned by the teacher from mixed-quality offline data into a single-step sampling model. The additional reward objective acts as a mode selection mechanism: effectively steering the student planner toward selecting high reward modes from the multi-modal distributions captured by the teacher diffusion planner.
@@
Our decoupled training framework offers two additional key advantages: training simplicity and flexibility. By generating clean action trajectories through single-step diffusion sampling, our approach removes the need for noise-aware reward models required for classifier-guided sampling from multi-step diffusion models (Janner et al., 2022), which also suffer from reduced accuracy at high noise levels. Furthermore, the reward model is trained independently from the teacher diffusion planner and the distillation process, avoiding the tight coupling that arises from concurrent multi-network training present in actor-critic methods. By integrating reward optimization directly into the single-step distillation process, our method simultaneously enables efficient generation and achieves superior performance, while maintaining straightforward training procedures.
@@
We demonstrate the performance and sampling efficiency of our RACTD on the suboptimal heterogeneous D4RL Gym-MuJoCo and D4RL FrankaKitchen benchmark and challenging long-horizon planning task Maze2d (Fu et al., 2020). Our method demonstrates both superior performance and substantial sampling efficiency compared to existing approaches, achieving an $9.7\%$ improvement compared to existing state-of-the-art (SOTA) while maintaining a $142$-fold reduction in sampling time inherited from the consistency trajectory distillation.
@@
Our contributions include: (1) We propose a novel reward-aware consistency trajectory distillation method for offline RL that enables single-step action trajectory generation while achieving superior performance, (2) We demonstrate that our approach enables decoupled training without the complexity of concurrent multi-network optimization or noise aware reward model training, (3) Through comprehensive experiments on multi-modal suboptimal dataset and long-horizon planning tasks, we show that our method achieves $9.7\%$ improvement over prior SOTA while achieving up to $142\times$ speedup by leveraging consistency trajectory distillation. We will publicly release our code upon acceptance of this paper.
@@ 2.3 Consistency Trajectory Distillation
Since the distance metric $d$ is defined on the clean data space and may not be well-defined in the noisy data space, in practice we further map all predictions to time 0 using the student model. Formally, denote the student model as $G_{\theta}$, the CTM loss is defined as:
$\mathcal{L}_{\mathrm{CTM}}=\mathbb{E}\left[d\left(G_{sg(\theta)}(\hat{\bm{x}}_{k}^{(t)},k,0),G_{sg(\theta)}(\bm{x}_{k}^{(t,u)},k,0)\right)\right]$ (5)
@@
In addition to the CTM loss, Kim et al. (2023) also incorporates the DSM loss to enforce the generated samples to be close to the training data. The DSM loss for the student model is the same as the one for EDM in Equation 4:
@@
CTM loss together with DSM loss constitutes consistency trajectory distillation (CTD), resembling the model in Kim et al. (2023) without the optimal GAN loss. In Figure 2 we provide a visualization of these objectives on a PFODE trajectory. After the distillation, the student model can perform “anytime-to-anytime” jumps along the PFODE trajectory. One-step sampling can then be achieved by calculating $\hat{\bm{x}}^{(T)}_{0}=G_{\theta}(\bm{x}_{T},T,0)$. This one-step sampling ability unlocks significant speedups and many potential design choices that are beneficial in RL, which we will elaborate next section.
@@ 3.1 Motivation and Intuition
Diffusion models and their consistency-based counterparts have demonstrated promising results in decision-making tasks, particularly in capturing multimodal behavior patterns (Janner et al., 2022; Chi et al., 2023; Ajay et al., 2022). Common recipes for using these models in decision-making generally fall into one of the three paradigms: (1) training a diffusion or consistency planner on expert demonstrations via behavior cloning and deploying it directly as a policy (Chi et al., 2023; Prasad et al., 2024); (2) integrating diffusion or consistency models into actor-critic frameworks for RL tasks (Wang et al., 2022; Hansen-Estruch et al., 2023; Ding & Jin, 2023); or (3) using reward agnostic diffusion planner with guided diffusion sampling for RL tasks (Janner et al., 2022; Ajay et al., 2022).
While behavior cloning pipelines perform well when trained on expert demonstrations, they often struggle with suboptimal datasets (e.g., medium-replay buffers) collected from diverse behavior policies. Such datasets typically exhibit complex multimodal behavior patterns, where only some modes lead to high rewards. Although one could potentially use rejection sampling to filter out low-reward training data, this approach becomes prohibitively sample inefficient, particularly as the quality of the training data deteriorates. On the other hand, to generate high reward actions, actor-critic approaches require concurrent training of multiple neural networks with sensitive hyperparameters. Finally, guided diffusion sampling necessitates training noise-aware reward models and multi-step sampling, which could be detrimental for time-sensitive decision-making tasks like self-driving.
So how can we better leverage potentially suboptimal datasets to design a diffusion-based single-step sampling model with a simple training procedure? Our key idea is to utilize the multimodal information captured by the teacher diffusion planner and encourage the student diffusion planner to favor the high reward modes. We achieve this by incorporating a reward objective directly in the consistency distillation process. Since our student model can achieve single-step denoising, we can incorporate a reward model trained in the clean sample space and avoid the multi-step reward optimization for diffusion models.
@@ 3.2 Modeling Action Sequences
When applying diffusion models to RL, several modeling choices are available: modeling actions (as a policy), modeling rollouts (as a planner), or modeling state transitions (as a world model). Following Chi et al. (2023), we adopt a planner approach and model a fixed-length sequence of future actions conditioned on a fixed-length sequence of observed states. This formulation ensures consecutive actions form coherent sequences, and reduces the chances of generating idle actions.
Formally, let $\vec{\bm{s}}_{n}=\{\bm{s}_{n-h},\bm{s}_{n-h+1},\dots,\bm{s}_{n}\}$ denote a length-$h$ sequence of past states at rollout time $n$, and $\vec{\bm{a}}_{n}=\{\bm{a}_{n},\bm{a}_{n+1},\dots,\bm{a}_{n+c}\}$ represent a length-$c$ sequence of future actions. Both the teacher and the student learn to model the conditional distribution $p(\vec{\bm{a}}_{n}\mid\vec{\bm{s}}_{n})$. In the context of diffusion notations, $\bm{x}=\vec{\bm{a}}_{n}\mid\vec{\bm{s}}_{n}$.
@@ 3.3 Reward-Aware Consistency Trajectory Distillation
Based on our action trajectory formulation, we now present our approach to integrating reward optimization into the consistency trajectory distillation process.
Let $R_{\psi}$ be a pre-trained differentiable return-to-go network (i.e. reward model) that takes the state $\bm{s}_{n}$ and action $\bm{a}_{n}$ at rollout time $n$ as inputs and predicts the future discounted cumulative reward $\hat{r}_{n}=\sum_{j=0}^{H-n}\gamma^{j}r_{n+j}$. We implement this reward model with four ConvBlocks used by previous reward-guided sampling method (Janner et al., 2022), followed by a Linear layer to map output to the correct dimension (Appendix B). When the student model generates a prediction $\vec{\bm{a}}_{n}\mid\vec{\bm{s}}_{n}=\hat{\bm{x}}^{(T)}_{0}=G_{\theta}(\bm{x}_{T},T,0)$, we extract the action at time $n$, denoted as $\hat{\bm{a}}_{n}$, from the predicted sequence and pass it along with $\bm{s}_{n}$ to the frozen reward model $R_{\psi}$ to estimate $\hat{r}_{n}$. The goal of our reward-aware training is to maximize the estimated discounted cumulative reward. Mathematically, the reward objective is defined as
@@
where $\alpha$, $\beta$, and $\sigma$ are hyperparameters to balance different loss terms. This objective bears resemblance to established offline RL algorithms like off-policy deterministic policy gradient (Silver et al., 2014) (Appendix G), suggesting our approach builds upon well recognized theoretical foundations.
@@ 3.4 Decoupled Training
A key advantage of our method of combining reinforcement learning, diffusion models, and consistency distillation is the ability to support fully decoupled training of all components. Traditional actor-critic frameworks, which are commonly used to incorporate diffusion models into reinforcement learning, require simultaneous training of both the actor and critic networks from scratch. This concurrent optimization presents considerable challenges, often demanding extensive hyperparameter tuning and careful balancing of different learning objectives.
Guided diffusion sampling, as proposed in Janner et al. (2022), offers an alternative approach by taking inspiration from classifier guided diffusion (Dhariwal & Nichol, 2021; Song et al., 2020b). However, since these classifiers (i.e. reward models) must evaluate noisy states at every denoising step to provide gradient guidance, they require noise-aware training and thus cannot be separately trained from the diffusion model. Also, predicting the correct reward from highly corrupted input could be very challenging, which can lead to inaccurate guidance that accumulates during its multi-step sampling process.
@@ 3.5 Reward Objective as Mode Selection
In offline RL, models often have to learn from datasets containing behaviors of varying quality. While diffusion models excel at capturing these diverse behavioral modes, they inherently lack the ability to differentiate between actions that lead to high versus low rewards. Our RACTD addresses this limitation by transforming the reward-agnostic teacher diffusion sampling distribution into one that preferentially samples from high-reward modes. We empirically verify this through a comparative analysis using the D4RL hopper-medium-expert dataset Fu et al. (2020), which contains an equal mixture of expert demonstrations and suboptimal rollouts from a partially trained policy.
@@
Figure 3 illustrates the reward distributions of rollouts sampled from three models: the unconditioned teacher, unconditioned student, and RACTD. The dataset (grey) exhibits two distinct modes corresponding to medium-quality and expert rollouts. The unconditioned teacher model (blue) accurately captures this bimodal distribution, and the unconditioned student model (orange) faithfully replicates it. In contrast, our RACTD (green) concentrates its samples on the higher-reward mode, demonstrating that our reward guidance effectively identifies and selects optimal behaviors from the teacher’s multi-modal distribution. We also include the discussion between sample diversity and mode selection in Appendix F.
@@ 4 Experiment
In this section, we conduct experiments to demonstrate: (1) the effectness of our RACTD in identifying high reward behavior patterns among multimodal mix quality data, (2) the expressiveness of our single step model compared to its multistep counterpart on complex high dimensional long horizon planning tasks, and (3) the speed-up achieved over the teacher model and existing policy-based diffusion models inherited from our appropriate application of CTD.
@@ 4.1 Offline RL
We compare our approach against a comprehensive set of baselines, including vanilla behavior cloning (BC) and Consistency Policy (Consistency BC) (Ding & Jin, 2023); model-free RL algorithms CQL (Kumar et al., 2020) and IQL (Kostrikov et al., 2021); model-based algorithms Trajectory Transformer (TT) (Janner et al., 2021), MOPO (Yu et al., 2020), MOReL (Kidambi et al., 2020), MBOP (Argenson & Dulac-Arnold, 2020); autoregressive model Decision Transformer (DT) Chen et al. (2021); diffusion-based planner Diffuser Janner et al. (2022); and diffusion/flow-based actor-critic methods Diffusion QL (Wang et al., 2022), Consistency AC (Ding & Jin, 2023), and Flow Q-learning (Park et al., 2025).
@@
We evaluate our method and the baselines on D4RL Gym-MuJoCo and FrankaKitchen benchmark (Fu et al., 2020). Gym-MuJoCo is a popular offline RL benchmark that contains three environments (hopper, walker2d, halfcheetah) of mixtures of varying quality data (medium-replay, medium, medium-expert). The FrankaKitchen dataset features a 9-DoF robotic arm performing multi-task manipulation across four kitchen objects (microwave, kettle, light, cabinets) with mixed-quality demonstrations (partial, mixed).
Following the setups in the prior works, we report the performance of both online and offline model selection if available in their original paper. Online model selection reports the best-performing checkpoint observed during training, while offline model selection reports the performance of the last training epoch. Results for non-diffusion-based models and Diffuser are taken from Janner et al. (2022), and results for Diffusion QL and Consistency AC/BC are sourced from their respective papers. The results for RACTD are reported as the mean and standard error over 100 planning seeds. We use past observation length $h=1$ and prediction horizon $c=16$ and closed-loop planning in all experiments following the same setup in previous diffusion planners (Chi et al., 2023; Prasad et al., 2024) and include the ablations on observation and planning horizons in Appendix E.6. The best score is emphasized in bold and the second-best is underlined.
Table 1: (Offline RL: Gym-MuJoCo) Performance and sampling efficiency (NFE: Number of Function Evaluations) of RACTD and a variety of baselines on the D4RL Gym-MuJoCo benchmark.
Medium Expert Medium Medium Replay Avg $\uparrow$ NFE $\downarrow$